# The Broader Landscape of Robustness in Algorithmic Statistics

### v1

@@ 2 Contamination
We will focus on an even stronger adversary, known as the $\eta$-strong contamination model. In this model, $n$ samples are first drawn i.i.d. from a distribution $\mathcal{D}$. An adversary is allowed to inspect these samples, and replace any $\eta n$ of them with arbitrary other datapoints. [^2] A common question is whether the algorithm has knowledge of the corruption fraction $\eta$. This turns out to not matter much [JOR22]. For simplicity, we will assume $\eta$ is known. The resulting (modified) dataset is provided to the algorithm.
This setting is more challenging than Huber’s contamination model, in that strong contamination allows the adversary to a) see the realization of the samples before choosing their contamination strategy (sometimes called an adaptive adversary); and b) modify points (or, equivalently, add and remove points), whereas a Huber adversary can only add points. [^3] Though it is beyond the scope of this piece, understanding how these differences in capability affect the power of an adversary remain at the frontier of research [BLMT22, BBKL23]. In particular, this captures the setting where samples are drawn from a distribution which is $\eta$-close to $\mathcal{D}$ in total variation distance, thus expressing situations such as model misspecification. Despite the power of the adversary in the strong contamination model, we will see that it is still possible to achieve compelling algorithmic results.
@@ 2.1 Univariate Contamination-Robust Estimation
Fortunately, there are contamination-robust estimators for the mean. The simplest example is the median: let our estimator $\hat{\mu}$ be the $50$th percentile of the dataset. It is not hard to argue that the previous single-point contamination scheme will not be able to shift the median of a Gaussian dataset that much, but the median also enjoys much stronger robustness properties. Its (folklore) guarantees are formalized in the following statement.
Proposition 2.1 (Corollary 1.15 of [DK22]).
@@
The rate enjoyed by the empirical median is the sum of two terms, a recurring pattern we will see for many robust estimators. The first term is the standard rate from the non-contamination-robust setting ($1/\sqrt{n}$, sometimes called the parametric rate), and we will refer to the excess term (in this case, $\eta$) as the cost of contamination. [^4] Throughout this article, we will de-emphasize focus on certain constant factors. For example, we will elide precise constant factors in the cost of contamination, though they are known in some cases (e.g., see [DKK+18] for analysis of the median). Similarly, the breakdown point of estimators (the maximum value of $\eta$ that can tolerated) enjoys significant study in the literature [Don82, ZJS20] but will not be a focus of our exposition. Such constants may be important practical considerations. It is not hard to show that this rate can not be improved – optimality of the parametric rate is folklore, and there exists an $\eta$-strong contamination of $\mathcal{N}(\mu,1)$ whose mean is at distance $\Omega(\eta)$, demonstrating optimality of the latter term.
@@ 2.2 Multivariate Contamination-Robust Estimation
Proposition 2.3 (Proposition 2.1 of [LRV16]).
Let $\mu\in\mathbb{R}^{d}$ and $\Sigma\in\mathbb{R}^{d\times d}$ be a diagonal covariance matrix with 00 in the first coordinate and $1$ in all other coordinates. There exists a distribution which is $\eta$-close to $\mathcal{N}(\mu,\Sigma)$ in total variation distance with geometric median $\hat{\mu}$ such that
@@ 2.3 Efficient Multivariate Contamination-Robust Estimation
Observe that a) this method is computable in polynomial time, and b) the cost of contamination is only $\eta\sqrt{\log 1/\eta}$, independent of the dimension $d$ [^7] There is evidence that this extra $\sqrt{\log 1/\eta}$ factor is unavoidable for computationally efficient estimators [DKS17].
Broadly speaking, computationally efficient contamination-robust estimators can be designed by first reasoning about properties of the dataset generated from a distribution, and then employing them to algorithmically limit the impact of contamination. We will explore contamination-robust Gaussian mean estimation through the lens of a notion of stability (following the presentation of [DK22], previously termed resilience by [SCV18]). We will first describe how Gaussian data enjoys stability of its empirical mean and covariance. We then show how to use this stability property to design an inefficient contamination-robust mean estimator. Finally, we discuss how to use these ideas to design a simple spectral method that can efficiently compute a contamination-robust estimate of the mean.
@@
Essentially, this lemma says that if a contamination of a stable dataset does not have large eigenvalues, then its empirical mean will be close to the true mean.
@@ 3 Sub-Gaussian Rates for Heavy-Tailed Mean Estimation
This proposition is again shown via a tail bound on the averaged distribution $\hat{\mu}$, which has variance at most $1/n$. The result follows from Chebyshev’s inequality, and this analysis is tight for the case when we have bounds on only the second moments [Cat12].
@@ 3.1 Sub-Gaussian Rates for Multivariate Data
In 2017, Lugosi and Mendelson [LM19] proposed a variant of what is now called a combinatorial center. [^10] To better capture the chronology of this rapidly evolving research area, we reference the original 2017 appearance on this result on arXiv, rather than [LM19]’s 2019 publication. This can be seen as a refinement of the simple median: it seeks to find a point which is close to a majority of the bucket means in every univariate projection. More formally, we have the following definition.
@@
The following key lemma, showing that the mean $\mu$ is a combinatorial center of the bucket means, is central to the approach introduced by Lugosi and Mendelson [LM19].
@@
A heavy-tailed dilemma:
@@ 3.2 Efficient Sub-Gaussian Rates for Multivariate Data
In 2018, [^11] Again, we reference the 2018 announcement of the result on arXiv, rather than the 2020 publication of [Hop20]. Hopkins [Hop20] managed to bypass these barriers, giving an efficient algorithm for computing a combinatorial center, and thus the first computationally efficient algorithm for mean estimation of heavy-tailed distributions with sub-Gaussian rates.
@@
$\max_{v\in\mathbb{R}^{d},b\in\mathbb{R}^{k}}$ $\sum_{i=1}^{k}b_{i}\text{ subject to}$ (QP-CC)
$b_{i}$$\in\{0,1\}$ $\forall i\in[k]$
@@
$b_{i}\sqrt{\lambda}$ $\leq\langle Y_{i}-\nu,b_{i}v\rangle$ $\forall i\in[k]$
@@
To interpret this quadratic program, the first constraint enforces that the $b_{i}$’s are indicators, and the second ensures that $v$ is a unit vector. The most important constraint is the third: if $b_{i}=1$, this “counts” (via the objective function) that $Y_{i}$ is distant from $\nu$ in the direction $v$ (otherwise, if $b_{i}=0$, the condition is vacuously satisfied). The optimization problem tries to find the direction $v$ (and the corresponding count) that maximizes the number of points that exceed this distance condition. If the worst case over all directions $v$ is at most $\gamma k$, then we certify that $\nu$ satisfies the definition of a $(\gamma,\lambda)$-combinatorial center. However, recall that this optimization problem is a quadratic program, which can not in general be solved efficiently.
Instead, Hopkins considers a semidefinite programming (SDP) relaxation of this optimization problem. Let $v\in\mathbb{R}^{d},V\in\mathbb{R}^{d\times d},b\in\mathbb{R}^{k},B\in\mathbb{R}^{k\times k},W\in\mathbb{R}^{k\times d}$ be the optimization variables. Then we can consider the following optimization problem.
$\max$ $\sum_{i=1}^{k}b_{i}\text{ subject to}$ (SDP-CC)
$B_{ii}$ $=b_{i}$ $\forall i\in[k]$
@@
$b_{i}\sqrt{\lambda}$ $\leq\langle Y_{i}-\nu,W_{i}\rangle$ $\forall i\in[k]$
@@
At this point, we know how to certify that the true mean $\mu$ is a combinatorial center. The question of how to find a combinatorial center remains. One approach, employed by Hopkins [Hop20], is based on the powerful Sum-of-Squares (SoS) proofs-to-algorithms framework [BS16]. This technique has enjoyed tremendous success in developing efficient algorithms for a range of problems within and beyond the realm of robustness [RSS18]. However, SoS-based algorithms can be technical and require significant background. In this article, we focus instead on a simpler and more intuitive descent-based approach subsequently introduced in 2019 by Cherapanamjeri, Flammarion, and Bartlett [CFB19].
Returning to the quadratic program for certification: suppose we try to certify a point $\nu$ which is not a combinatorial center. What would a solution to the quadratic program (QP-CC) look like? In addition to returning the indicator $b_{i}$’s (telling us which points are “far”), it produces a vector $v$, pointing in the direction of these far points. Intuitively, taking a step in this direction to $\nu+\zeta v$ (where $\zeta$ is some judiciously chosen step size) should give us a point which is “more central” to the dataset. We can iterate this process on $\nu+\zeta v$, repeatedly obtaining a new direction to step in and eventually finding and certifying a point $\hat{\mu}$ to be a combinatorial center.
Of course, this discussion revolves around obtaining solutions to the quadratic program (QP-CC), which we can’t do in polynomial time. It turns out that the same general approach still works even on the relaxed SDP (SDP-CC). Indeed, while initial works required a non-trivial rounding scheme to extract a direction from the matrix $V$, it was subsequently shown by Hopkins, Kamath, and Majid [HKM22] that one can use the direction $v$ in the SDP without further modification, despite the loss of the interpretation in the quadratic program described above.
@@ 3.3 Connections with Contamination
In fact, though not a focus of the original works, it is now considered folklore that with a careful setting of parameters, the combinatorial center in the median-of-means paradigm is automatically contamination robust, thus giving both types of robustness simultaneously! To see this, recall that our goal was to find a $(0.01,O(r^{2}_{\beta}))$-combinatorial center of the bucket means. This notion is naturally contamination robust: if (say) $0.001k$ of the bucket means were modified, then the “centrality” (i.e., the score function of (QP-CC)) of any candidate center $\nu$ changes by at most $0.001k$. Therefore, if we find a $(0.01,O(r^{2}_{\beta}))$-combinatorial center under such contamination, it was (at worst) a $(0.011,O(r^{2}_{\beta}))$-combinatorial center prior to contamination. Such loosening of the parameters turns out to be OK throughout the analysis, in particular for the relaxed SDP (SDP-CC).
@@
Others have subsequently investigated algorithmic connections between contamination and sub-Gaussian rates [DL22, PBR19]. In a 2020 work, Hopkins, Li, and Zhang [HLZ20] made an important conceptual contribution, relating the two solution concepts we have seen thus far. Specifically, they showed that the spectral center (Definition 2.7) and combinatorial center (Definition 3.6) are equivalent to each other, up to some gap in constants (see Propositions 5.1 and 5.2 of [HLZ20]). Algorithmically, this shows that one can use either the combinatorial or the spectral center with the median-of-means paradigm to achieve sub-Gaussian rates. In a similar vein, Diakonikolas, Kane, and Pensia [DKP20] simultaneously showed that, more generally, any contamination-robust estimator based on stability (of the sort used in Definition 2.5) can be employed as an aggregator in the median-of-means paradigm to achieve the sub-Gaussian rate.
@@ 4 Differential Privacy
An algorithm $A\ :\ \mathcal{X}^{n}\rightarrow\mathcal{Y}$ is $(\varepsilon,\delta)$-differentially private (DP) if, for all datasets $X,X^{\prime}\in\mathcal{X}^{n}$ that differ in one entry and for all $S\subseteq\mathcal{Y}$, we have that
@@ 4.1 Privacy Fundamentals and Three Mean Estimators
We will focus on private mean estimation under differential privacy, restricting our attention to distributions $\mathcal{D}$ with bounded covariance $\Sigma\preceq I$ [^13] There is also significant study into private Gaussian mean estimation (see, e.g., [KV18, KLSU19, BKSW19, BGS+21, BHS23, KDH23, HKMN23]), but we focus on the bounded covariance setting which captures the most interesting difficulties. Given a dataset $X_{1},\dots,X_{n}\in\mathbb{R}^{d}$, our goals for the estimator $\hat{\mu}(X_{1},\dots,X_{n})$ are twofold:
@@
Let $X_{1},\dots,X_{n}$ be $n$ samples from a distribution $\mathcal{D}$ with mean $\mu\in\mathbb{R}^{d}$ such that $\|\mu\|_{2}\leq\sqrt{d}$ and covariance $\Sigma\preceq I$ [^15] In the differential private setting, many estimators incur some dependence on a bound $R$ for the mean, i.e., a value such that $\|\mu\|_{2}\leq R$. In many cases this cost is logarithmic in $R$, and is shown to be necessary by matching lower bounds. To streamline our presentation, we focus on the case where we start with a “coarse estimate” for the mean $\mu$ with $\ell_{2}$-error $\sqrt{d}$ and re-center around it, yielding the condition $\|\mu\|_{2}\leq\sqrt{d}$. There exists a polynomial-time $(\varepsilon,0)$-differentially private estimator $\hat{\mu}(X_{1},\dots,X_{n})$ such that, with probability $\geq 95\%$,
@@
Let $\mathcal{X}^{n}$ be the set of $n$ element datasets, $\mathcal{H}$ be a set of objects, and $s\ :\ \mathcal{X}^{n}\times\mathcal{H}\rightarrow\mathbb{R}$ be a score function that outputs the “quality” of an object $h$ with respect to a dataset $X$. Let
@@
While the description of the exponential mechanism is somewhat abstract, it is quite a general and expressive primitive. For example, the Laplace mechanism can be phrased as a special case: for a univariate function $f$, one can instantiate the exponential mechanism with $\mathcal{H}=\mathbb{R}$ and $s(X,h)=-|f(X)-h|$.
@@
To instantiate the exponential mechanism, we need to define a set of objects and a score function. Our objects will be candidates for the estimate of the mean: that is, a subset of $\Theta=\{\nu\in\mathbb{R}^{d}\ :\ \|\nu\|_{2}\leq\sqrt{d}\}$. To make our set of objects finite, we consider a cover of $\Theta$, a subset $\mathcal{H}\subseteq\Theta$ such that for every $\nu\in\Theta$, there exists a nearby $\nu^{\prime}\in\mathcal{H}$ such that $\|\nu-\nu^{\prime}\|_{2}$ is small. Note that the size of such a cover is necessarily exponential in the dimension $d$. Thus, naively, the running time of the exponential will also be exponential in $d$. On the other hand, observe that the loss in utility (described by Proposition 4.5) has only a mild logarithmic dependence on the size of $\mathcal{H}$ (and thus a linear dependence on the dimension $d$) – this will be important when bounding the final error.
@@
The more substantial difference in our application of (QP-CC) (in comparison to that in Section 3) is the number of pieces $k$ we partition our data into. Recall that previously, we set $k=\Omega(\log 1/\beta)$ to achieve sub-Gaussian rates, and $k=\Omega(\eta n)$ to achieve robustness. This time, we will set $k=\tilde{\Omega}(d/\varepsilon)$ for privacy. Again, we split the dataset into $k$ parts and compute the bucket means of each, which we denote as $Y_{1},\dots,Y_{k}$. Following Lemma 3.7, we can see that the mean $\mu$ (and, to account for the discretization of the cover, any point which is sufficiently close) is a $\left(0.01,\tilde{O}\left(\frac{d}{\varepsilon n}\right)\right)$-combinatorial center. This implies that the objective function of (QP-CC) for such a point will be $\geq 0.99k=\tilde{\Omega}(d/\varepsilon)$, and, by the utility guarantees of Proposition 4.5, the loss of score due to running the exponential mechanism will be only $O\left(\frac{\Delta}{\varepsilon}\log|\mathcal{H}|\right)=\tilde{O}\left(d/\varepsilon\right)$. This implies that the exponential mechanism will still return a combinatorial center of the dataset, and consequently, a low-error estimate of the mean as described in Theorem 4.6. The algorithm is describe more precisely in Figure 2.

### v2

@@ 1 Introduction
Table 1: Summary of different types of robustness considered in this article.
Type of Robustness Input data Estimator’s Goal Section
Contamination Stochastic, a fraction is modified arbitrarily Bounded error, depending on corruption rate Section 2
Heavy-Tailed Data Stochastic, from a distribution with heavy tails Mild dependence of error on (inverse) failure probability Section 3
Privacy Arbitrary, a single point is modified arbitrarily Approximately preserve estimator’s output distribution Section 4
@@ 2 Contamination
We will focus on an even stronger adversary, known as the $\eta$-strong contamination model. In this model, $n$ samples are first drawn i.i.d. from a distribution $\mathcal{D}$. An adversary is allowed to inspect these samples, and replace any $\eta n$ of them with arbitrary other datapoints. [^3] A common question is whether the algorithm has knowledge of the corruption fraction $\eta$. In most cases, this turns out to not matter much [JOR22] (though there are exceptions, e.g., for identifiability reasons when both $\eta$ and the covariance matrix are unknown, see Exercise 1.7 (b) of [DK19]). For simplicity, we will assume $\eta$ is known. The resulting (modified) dataset is provided to the algorithm.
This setting is more challenging than Huber’s contamination model, in that strong contamination allows the adversary to a) see the realization of the samples before choosing their contamination strategy (sometimes called an adaptive adversary); and b) modify points (or, equivalently, add and remove points), whereas a Huber adversary can only add points. [^4] Though it is beyond the scope of this piece, understanding how these differences in capability affect the power of an adversary remain at the frontier of research [BLMT22, BBKL23, BV25, LBK25]. In particular, this captures the setting where samples are drawn from a distribution which is $\eta$-close to $\mathcal{D}$ in total variation distance, thus expressing situations such as model misspecification. Despite the power of the adversary in the strong contamination model, we will see that it is still possible to achieve compelling algorithmic results.
@@ 2.1 Univariate Contamination-Robust Estimation
Fortunately, there are contamination-robust estimators for the mean. The simplest example is the median: let our estimator $\hat{\mu}$ be the $50$th percentile of the dataset. It is not hard to argue that the previous single-point contamination scheme will not be able to shift the median of a Gaussian dataset that much, but the median also enjoys much stronger robustness properties [HR09]. Its (folklore) guarantees are formalized in the following statement.
Proposition 2.1 (e.g., Corollary 1.15 of [DK22]).
@@
The rate enjoyed by the empirical median is the sum of two terms, a recurring pattern we will see for many robust estimators. The first term is the standard rate from the non-contamination-robust setting ($1/\sqrt{n}$, sometimes called the parametric rate), and we will refer to the excess term (in this case, $\eta$) as the cost of contamination. [^5] Throughout this article, we will de-emphasize focus on certain constant factors. For example, we will elide precise constant factors in the cost of contamination, though they are known in some cases (e.g., see [DKK+18] for analysis of the median). Similarly, the breakdown point of estimators (the maximum value of $\eta$ that can tolerated) enjoys significant study in the literature [Don82, ZJS20, CSS25] but will not be a focus of our exposition. Such constants may be important in practical considerations. It is not hard to show that this rate can not be improved – optimality of the parametric rate is textbook (see, e.g., Example 15.4 of [Wai19]), and there exists an $\eta$-strong contamination of $\mathcal{N}(\mu,1)$ whose mean is at distance $\Omega(\eta)$, demonstrating optimality of the latter term.
@@ 2.2 Multivariate Contamination-Robust Estimation
Proposition 2.3 (Proposition 2.1 of [LRV16], see also Lemma 1 of [PBR20] for a more general statement).
Let $\mu\in\mathbb{R}^{d}$ and $\Sigma\in\mathbb{R}^{d\times d}$ be a diagonal covariance matrix with 0 in the first coordinate and $1$ in all other coordinates. There exists a distribution which is $\eta$-close to $\mathcal{N}(\mu,\Sigma)$ in total variation distance with geometric median $\hat{\mu}$ such that
@@ 2.3 Efficient Multivariate Contamination-Robust Estimation
Observe that a) this method is computable in polynomial time, and b) the cost of contamination is only $\eta\sqrt{\log 1/\eta}$, independent of the dimension dd. [^8] There is evidence that this extra $\sqrt{\log 1/\eta}$ factor is unavoidable for computationally efficient estimators [DKS17].
Broadly speaking, computationally efficient contamination-robust estimators can be designed by first reasoning about properties of the dataset generated from a distribution, and then employing them to algorithmically limit the impact of contamination. We will explore contamination-robust Gaussian mean estimation through the lens of a notion of stability (following the presentation of [DK22], note also the closely related concept of resilience [SCV18]). We will first describe how Gaussian data enjoys stability of its empirical mean and covariance. This will allow us to expose the core solution concepts that underlie the resulting algorithms, and set things up to subsequently discuss connections with other forms of robustness. We then show how to use these properties to design an inefficient contamination-robust mean estimator. Finally, we discuss a simple spectral method based on these ideas, that can efficiently compute a contamination-robust estimate of the mean.
@@
Essentially, this lemma says that if a contamination of a stable dataset does not have large variance in any direction (i.e., the eigenvalues of its empirical covariance matrix are not too large), then its empirical mean will be close to the true mean.
@@ 3 Sub-Gaussian Rates for Heavy-Tailed Mean Estimation
Furthermore, there exists such a distribution $\mathcal{D}$ such that, with probability $\geq 1-\beta$,
$|\hat{\mu}-\mu|\geq\Omega\left(\sqrt{\frac{1}{\beta n}}\right).$
The upper bound in this proposition is again shown via a tail bound on the averaged distribution $\hat{\mu}$, which has variance at most $1/n$. The result follows from Chebyshev’s inequality. This analysis is tight for the case when we have bounds on only the second moments [Cat12].
@@ 3.1 Sub-Gaussian Rates for Multivariate Data
In 2017, Lugosi and Mendelson [LM17, LM19] proposed a variant of what is now called a combinatorial center. This can be seen as a refinement of the simple median: it seeks to find a point which is close to a majority of the bucket means in every univariate projection. More formally, we have the following definition.
@@
The combinatorial center is an essential concept in many recent algorithms achieving sub-Gaussian rates. The following key lemma, showing that the mean $\mu$ is a combinatorial center of the bucket means, is central to the approach introduced by Lugosi and Mendelson [LM19].
@@
A heavy-tailed dilemma
@@ 3.2 Efficient Sub-Gaussian Rates for Multivariate Data
In 2018, Hopkins [Hop18, Hop20] managed to bypass these barriers, giving an efficient algorithm for computing a combinatorial center, and thus the first computationally efficient algorithm for mean estimation of heavy-tailed distributions with sub-Gaussian rates. We will start by describing an inefficient approach to find a combinatorial center, and then discuss ideas that allow us to make it efficient.
@@
$\max_{v\in\mathbb{R}^{d},b\in\mathbb{R}^{k}}$ $\sum_{i=1}^{k}b_{i}\text{ subject to}$ (QP-$C$)
$b_{i}$ $\in\{0,1\}\$$\forall i\in[k]$
@@
$b_{i}\sqrt{\lambda}$ $\leq\langle Y_{i}-\nu,b_{i}v\rangle\$$\forall i\in[k]$
@@
To interpret this quadratic program, the first constraint enforces that the $b_{i}$’s are indicators, and the second ensures that $v$ is a unit vector. The most important constraint is the third: if $b_{i}=1$, this “counts” (via the objective function) that $Y_{i}$ is distant from $\nu$ in the direction $v$ (otherwise, if $b_{i}=0$, the condition is vacuously satisfied). The optimization problem tries to find the direction $v$ (and the corresponding count) that maximizes the number of points that exceed this distance condition. If the worst case over all directions $v$ is at most $\gamma k$, then we certify that $\nu$ satisfies the definition of a $(\gamma,\lambda)$-combinatorial center.
Given this procedure that certifies that a point is a combinatorial center, how do we find a combinatorial center? If we are not concerned with computation, a straightforward approach is to discretize the space and try every point $\nu$ as a candidate combinatorial center in the quadratic program (QP-CC). However, we will instead employ an oracle-efficient descent-based approach introduced in 2019 by Cherapanamjeri, Flammarion, and Bartlett [CFB19]. This has the advantage that, given an oracle that solves the quadratic program (QP-CC), one can find a combinatorial center with only $O(\log d)$ calls to the oracle. This leaves us with only one source of computational intractability (solving the quadratic program) rather than introducing a new one.
Suppose we tried to certify a point $\nu$ which is not a combinatorial center. What would a solution to the quadratic program (QP-CC) look like? In addition to returning the indicator $b_{i}$’s (telling us which points are “far”), it produces a vector $v$, pointing in the direction of these far points. Intuitively, taking a step in this direction to $\nu+\zeta v$ (where $\zeta$ is some judiciously chosen step size) should give us a point which is “more central” to the dataset. We can iterate this process on $\nu+\zeta v$, repeatedly obtaining a new direction to step in and eventually finding and certifying a point $\hat{\mu}$ to be a combinatorial center.
Now that we have an inefficient algorithm, we turn our attention to making the procedure efficient. At the heart of this algorithm, the intractability arises due to the quadratic program for certification (QP-CC). The canonical way to make an optimization problem based on quadratic programming efficient is by instead considering its semidefinite programming (SDP) relaxation. Let $v\in\mathbb{R}^{d},V\in\mathbb{R}^{d\times d},b\in\mathbb{R}^{k},B\in\mathbb{R}^{k\times k},W\in\mathbb{R}^{k\times d}$ be the optimization variables. Hopkins [Hop20] introduced the following optimization problem.
$\max$ $\sum_{i=1}^{k}b_{i}\text{ subject to}$ (SDP-$C$)
$B_{ii}$ $=b_{i}\$$\forall i\in[k]$
@@
$b_{i}\sqrt{\lambda}$ $\leq\langle Y_{i}-\nu,W_{i}\rangle\$$\forall i\in[k]$
@@
This establishes how to efficiently solve the certification problem. Hopkins [Hop20] goes on to use the powerful Sum-of-Squares (SoS) proofs-to-algorithms framework to find a (certifiable) combinatorial center. [^11] See the textbook of Barak and Steurer for an introduction to SoS [BS16]. This technique has enjoyed tremendous success in developing efficient algorithms for a range of problems within and beyond the realm of robustness [RSS18]. However, SoS-based algorithms can be technical and require significant background.
It turns out the aforementioned descent-based approach of Cherapanamjeri, Flammarion, and Bartlett [CFB19] still works even on the relaxed SDP (SDP-CC). Indeed, while initial works required a non-trivial rounding scheme to extract a direction from the matrix $V$, it was subsequently shown by Hopkins, Kamath, and Majid [HKM22] that one can use the direction $v$ in the SDP without further modification, despite the loss of the interpretation in the quadratic program described above.
@@ 3.3 Connections with Contamination
In fact, though not a focus of the original works (and first observed by [DL22]), it is now well understood that with a careful setting of parameters, the combinatorial center in the median-of-means paradigm is automatically contamination robust, thus giving both types of robustness simultaneously! To see this, recall that our goal was to find a $(0.01,O(r^{2}_{\beta}))$-combinatorial center of the bucket means. This notion is naturally contamination robust: if (say) $0.001k$ of the bucket means were modified, then the “centrality” (i.e., the score function of (QP-CC)) of any candidate center $\nu$ changes by at most $0.001k$. Therefore, if we find a $(0.01,O(r^{2}_{\beta}))$-combinatorial center under such contamination, it was (at worst) a $(0.011,O(r^{2}_{\beta}))$-combinatorial center prior to contamination. Such loosening of the parameters turns out to be OK throughout the analysis, in particular for the relaxed SDP (SDP-CC).
@@
Others have subsequently investigated algorithmic connections between contamination and sub-Gaussian rates [DL22, PBR20]. In a 2020 work, Hopkins, Li, and Zhang [HLZ20] made an important conceptual contribution, relating the two solution concepts we have seen thus far. Specifically, they showed that the spectral center (Definition 2.7) and combinatorial center (Definition 3.6) are equivalent to each other, up to some gap in constants (see Propositions 5.1 and 5.2 of [HLZ20]). Algorithmically, this shows that one can use either the combinatorial or the spectral center with the median-of-means paradigm to achieve sub-Gaussian rates. Simultaneously, Diakonikolas, Kane, and Pensia [DKP20] showed the more general result that any contamination-robust estimator based on stability (of the sort used in Definition 2.5) can be employed as an aggregator in the median-of-means paradigm to achieve the sub-Gaussian rate. In fact, they further show that such estimators can be used directly to achieve a near sub-Gaussian rate, entirely bypassing the median-of-means approach. While this fact is conceptually interesting in its own right, it also allows one to employ higher-order moment information when establishing stability, leading to better contamination robustness than the $\sqrt{\eta}$ of Theorem 3.12.
@@ 4 Differential Privacy
[…]
@@ 4.1 Privacy Fundamentals and Three Mean Estimators
We will focus on private mean estimation under differential privacy, restricting our attention to distributions $\mathcal{D}$ with bounded covariance $\Sigma\preceq I$. [^13] There is also significant study into private Gaussian mean estimation (see, e.g., [KV18, KLSU19, BKSW19, BGS+21, BHS23, KDH23, HKMN23]), but we focus on the bounded covariance setting which captures the most interesting difficulties. Given a dataset $X_{1},\dots,X_{n}\in\mathbb{R}^{d}$, our goals for the estimator $\hat{\mu}(X_{1},\dots,X_{n})$ are twofold:
@@
We will first introduce three algorithms for private mean estimation, each deficient in a different way, and then describe an algorithm which overcomes these issues. The reader solely interested in connections with other forms of robustness may safely skip the first two.
@@
Let $X_{1},\dots,X_{n}$ be $n$ samples from a distribution $\mathcal{D}$ with mean $\mu\in\mathbb{R}^{d}$ such that $\|\mu\|_{2}\leq\sqrt{d}$ and covariance $\Sigma\preceq I$. [^15] In the differential private setting, many estimators incur some dependence on a bound $R$ for the mean, i.e., a value such that $\|\mu\|_{2}\leq R$. In many cases this cost is logarithmic in $R$, and is shown to be necessary by matching lower bounds. To streamline our presentation, we focus on the case where we start with a “coarse estimate” for the mean $\mu$ with $\ell_{2}$-error $\sqrt{d}$ and re-center around it, yielding the condition $\|\mu\|_{2}\leq\sqrt{d}$. There exists a polynomial-time $(\varepsilon,0)$-differentially private estimator $\hat{\mu}(X_{1},\dots,X_{n})$ such that, with probability $\geq 95\%$,
@@
[…]
@@
While the description of the exponential mechanism is somewhat abstract, it is quite a general and expressive primitive. For example, the Laplace mechanism can be phrased as a special case: for a univariate function $f$, one can instantiate the exponential mechanism with $\mathcal{H}=\mathbb{R}$ and s(X,h)=|f(X)h|s(X,h)=-|f(X)-h|.
@@
To instantiate the exponential mechanism, we need to define a set of objects and a score function. Our objects will be candidates for the estimate of the mean: that is, a subset of $\Theta=\{\nu\in\mathbb{R}^{d}\ :\ \|\nu\|_{2}\leq\sqrt{d}\}$. To make our set of objects finite, we consider a cover of $\Theta$, a subset $\mathcal{H}\subseteq\Theta$ such that for every $\nu\in\Theta$, there exists a nearby $\nu^{\prime}\in\mathcal{H}$ such that $\|\nu-\nu^{\prime}\|_{2}$ is small. Note that the size of such a cover is necessarily exponential in the dimension $d$. Thus, naively, the running time of the exponential mechanism will also be exponential in $d$. On the other hand, observe that the loss in utility (described by Proposition 4.5) has only a mild logarithmic dependence on the size of $\mathcal{H}$ (and thus a linear dependence on the dimension $d$) – this will be important when bounding the final error.
@@
The more substantial difference in our application of (QP-CC) (in comparison to that in Section 3) is the number of pieces $k$ we partition our data into. Recall that previously, we set $k=\Omega(\log 1/\beta)$ to achieve sub-Gaussian rates, and $k=\Omega(\eta n)$ to achieve robustness. This time, we will set $k=\tilde{\Omega}(d/\varepsilon)$ for privacy. Again, we split the dataset into $k$ parts and compute the bucket means of each, which we denote as $Y_{1},\dots,Y_{k}$. Following Lemma 3.7, we can see that the mean $\mu$ (and, to account for the discretization of the cover, any point which is sufficiently close) is a $\left(0.01,\tilde{O}\left(\frac{d}{\varepsilon n}\right)\right)$-combinatorial center with high probability. This implies that the objective function of (QP-CC) for such a point will be $\geq 0.99k=\tilde{\Omega}(d/\varepsilon)$, and, by the utility guarantees of Proposition 4.5, the loss of score due to running the exponential mechanism will be only $O\left(\frac{\Delta}{\varepsilon}\log|\mathcal{H}|\right)=\tilde{O}\left(d/\varepsilon\right)$. This implies that the exponential mechanism will still return a combinatorial center of the dataset, and consequently, a low-error estimate of the mean as described in Theorem 4.6. The algorithm is describe more precisely in Figure 2.